OpenStudy (anonymous):

find an equation in standard form thats passing through the origin and with slope 0

OpenStudy (anonymous):

y=0

OpenStudy (anonymous):

in standard form , equation ?

OpenStudy (anonymous):

or y=0x+b.... b equaing 0 to pass through the origin

OpenStudy (jdoe0001):

well, let's use the point-slope form, see what we get \(\large y-y_1=m(x-x_1)\)

OpenStudy (jdoe0001):

so a point is (0, 0), slope = m = 0 y-0 = 0(x-0)